1/1

N. N. 43,  N. N. 43,
Ein irischer Weihnachts-Segenswunsch  An Irish Christmas Blessing

tradukita de Gudrun Hickey

Gott gewähre Dir die Leichtigkeit in Deinen Schritten, Ein Lächeln in jedem Gesicht, auf das Du triffst, Liebe Menschen an Deinem Ofen versammelt, Und gute Freunde an Deiner Tür, Ein Weihnachtslied auf Deinen Lippne, Eine Kerze, die hell am Fenster brennt, Und möge der Herr Dein Herz beglücken, Und zur Heiligen Nacht in ihm verweilen.

God grant you lightness in your step,

a smile on every face you meet, Loved ones gathered at your hearth, and at your door, good friends to greet A holy hymn upon your lips, a window candle burning bright, And may the Good Lord bless your Heart and come to dwell there at Christmas night.

Traduko de la Angla poemo “An Irish Christmas Blessing” de N. N. 43 en la Germanan de Gudrun Hickey.

Arg-723-1457 (2012-12-13 21:58:50)

Verkinto de tiu ĉi Angla poemo estas N. N. 43.

Arg-723-1456 (2012-12-13 22:05:43)

This Irish Christmas Blessing can be found in http: // www. irishcultureandcustoms. com/ Blessings/ Bless2. html .

http://www.poezio.net/version?poem-id=723&version-id=1457,1456